OpenStudy (anonymous):

Jane had a bottle filled with juice. At first, she drank 1/5 of the juice in the bottle. After 1 hour, she drank 1/4 of the juice remaining in the bottle. After another 2 hours, she drank 1/3 of the remaining juice in the bottle. At that point, Jane checked how much juice was left: There was 2/3 of a cup left. How much juice was in the bottle originally?

OpenStudy (campbell_st):

ok... so she drank 1/5 so 4/5 remaining... so 1/4 of 4/5 = 1/5 so 3/5 remaining then 1/3 of 3/5 = 1/5 so there is 3/5 been drunk and 2/5 remaining.... so 2/5 = 2/3 of a cup... or dividing by 2 to find the measure of 1/5 1/5 = 1/3 of a cup then 5 lots of 1/5 = 5 * 1/3 the the drink contained 5/3 or 1 and 2/3 cups hope it makes sense

OpenStudy (anonymous):

Oh! I totally get what you're saying. Thanks so much, this is really helpful and I appreciate it a lot! :)